OpenStudy (anonymous):

for a fan and medal The midpoint of PQ is (–6, –3). One endpoint of PQ is P(–1, –4). What are the coordinates of Q? A. Q(–11, –2) B. Q(–3.5, –3.5) C. Q(2, 11) D. Q(11, 2)

OpenStudy (praxer):

not the section formula use ( x2+x1)/2 =X and Y= ( y2+y1)/2

OpenStudy (anonymous):

alright

OpenStudy (praxer):

check you answer, i think it is a as -11+(-1) = -12/2 = -6.. check it out, once again.

OpenStudy (anonymous):

yep i got it right

OpenStudy (praxer):

:)

OpenStudy (anonymous):

thanks!!

OpenStudy (praxer):

welcome... :)